OpenStudy (anonymous):

i need to: Find f^-1(x) if f(x) =3sin(2x)cos(x)

OpenStudy (anonymous):

There's no inverse

OpenStudy (anonymous):

x = 3 (2 sin(x) cos(y)) cos(y), and express in terms of cos(y) only. You will then have a cubic equation in sin(y): \[\sin ^{3}x - \sin x + y/6 = 0,\] written in standard form. Solve for sin y, then take the arcsine of the result. The exact solution will be complicated!

OpenStudy (anonymous):

It is \[ f ^{-1}(x) = \sin ^{-1}(\sqrt[3]{-y/12 + (1/2) \sqrt{-4 +y ^{2}/36}}\] \[+\sqrt[3]{-y/12- (1/2)\sqrt{-4+y ^{2}/36}}).\]

OpenStudy (anonymous):

We need to determine its domain.

OpenStudy (anonymous):

BTW, those y's are supposed to be x's.

OpenStudy (anonymous):

I though the original function failed the horizontal line test

OpenStudy (anonymous):

How can the inverse exist if that's the case, lol

OpenStudy (anonymous):

You need to restrict the domain of the original function for it to have a valid inverse.

OpenStudy (anonymous):

It will be a very small domain

OpenStudy (anonymous):

Graph the original function and look for intervals where the function is strictly increasing and where it's strictly decreasing.

OpenStudy (anonymous):

like not enough for it to be even considered the graph of the original function, I'm sorry, I'm going with no inverse

OpenStudy (anonymous):

You'd have to say that one domain is from -.615 to .615 and that's just one domain, you have to list all of them

OpenStudy (anonymous):

good luck with that

OpenStudy (anonymous):

I agree that the problem is quite complicated. It is what it is, and not all problems will be easy.

OpenStudy (anonymous):

Here's another wrinkle: Since the equation is cubic in sin x, there are two other roots. It would be nice if they were complex so we could disregard them. Are they?

OpenStudy (anonymous):

There is a lot to this problem.

OpenStudy (anonymous):

I think the question was meant to be graphed first and checked to see if it passes the horizontal line test

OpenStudy (anonymous):

If so, then I believe we are to consider all real numbers to be the domain of the original function.

OpenStudy (anonymous):

You would then be correct.

OpenStudy (anonymous):

I don't see why we wouldn't....It didn't give us a restricted domain to begin with.

OpenStudy (anonymous):

True.